OpenStudy (anonymous):

find the volume of a cone with a radius of 4 cm and a height of 10 cm

OpenStudy (anonymous):

The formula for finding the volume of a cylinder is \[V=Bh\] B being the base. You'll start off by finding the area of the base with the formula \[\pi r^{2}\] then multiply by the height.

OpenStudy (anonymous):

i got 150.24

ganeshie8 (ganeshie8):

doesnt look correct

OpenStudy (anonymous):

its not

ganeshie8 (ganeshie8):

cuz, r^2*h itself comes to 16*10 = 160

ganeshie8 (ganeshie8):

if u multiply it wid pi, it wud be atleast 3 times more than 160

ganeshie8 (ganeshie8):

check ur calculation once

OpenStudy (anonymous):

i got 1507.2

ganeshie8 (ganeshie8):

hey hang on, this is a cone ! not a cylinder

OpenStudy (anonymous):

yea

ganeshie8 (ganeshie8):

ok, then use this formula :- volume of cone = \(\frac{1}{3} \pi r^2 h\)

OpenStudy (anonymous):

ok i got 167.47

ganeshie8 (ganeshie8):

thats right !

OpenStudy (anonymous):

Thank You

ganeshie8 (ganeshie8):

np :) look at dan's its interesting integral for volume..

OpenStudy (dan815):

dont worry about my way its prolly ridiculously wrong! iam practicing

OpenStudy (dan815):

:)

OpenStudy (anonymous):

:) its ok

ganeshie8 (ganeshie8):

its advanced it looks beautiful :)

OpenStudy (dan815):

hhahah i wanted to try something

OpenStudy (dan815):